OpenStudy (anonymous):

what is the inverse equations of the following: a) y=sqrt(x-3)-1 b)y=(x-1)^2+3 c)y=2/x+5-1

OpenStudy (johnweldon1993):

Alright...so for the inverse equation...we just switch x and y... and solve for 'y' again... for the first one... \[\large y = \sqrt{x - 3} - 1\] becomes \[\large x = \sqrt{y - 3} - 1\] Now add 1 to both sides... \[\large x + 1 = \sqrt{y - 3}\] Now you square both sides to get rid of that radical... \[\large (x + 1)^2 = y - 3\] And finally you add 3 to both sides... \[\large y = (x + 1)^2 + 3\] And that is your inverse...can you do the other 2?